OpenStudy (rockinhood):

Will Fan and medal immediately for this easy question How do I Write these two phrases? x is equal to more than one but less than seven x is equal to more than four but less than fifteen

OpenStudy (danjs):

1<x<7

OpenStudy (danjs):

4<x<15

OpenStudy (anonymous):

x=1-7 x=4-15 I think ???

OpenStudy (ericwise):

DanJS is correct: 1<x<7 4<x<15 In worded form: 1 (is less than) x (is less than) 7 4 (is less than) x (is less than) 15

OpenStudy (ericwise):

does that makes sense?